If u=x(1-y) and v=xy, then find the value of the Jacobian ∂u,v∂(x,y)



Select one:



A -x




B x2




C -x2




D x


1
Expert's answer
2022-03-29T08:51:56-0400

"\\frac{\\partial \\left( u,v \\right)}{\\partial \\left( x,y \\right)}=\\left| \\begin{matrix}\t\\frac{\\partial u}{\\partial x}&\t\t\\frac{\\partial v}{\\partial x}\\\\\t\\frac{\\partial u}{\\partial y}&\t\t\\frac{\\partial u}{\\partial y}\\\\\\end{matrix} \\right|=\\left| \\begin{matrix}\t1-y&\t\ty\\\\\t-x&\t\tx\\\\\\end{matrix} \\right|=x\\left( 1-y \\right) +xy=x"